# Fulde-Ferrell-Larkin-Ovchinnikov States and Topological Bogoliubov Fermi Surfaces in Altermagnets: an Analytical Study

### v1

@@ III.2 Thermodynamic potential and ground state energy
In the ground state at zero temperature, referred to as the ground-state energy, $\Omega$ takes the following form,
@@
where $\Theta(x)$ is the Heaviside step function. Converting the momentum sum to an integration, the ground-state energy density becomes:
@@
where the renormalization condition Eq. (4) is used. At a given chemical potential, the phase boundary between SF and N phase is obtained by equating the ground-state energies of these two phases:
@@ III.3 Gap equation
During the phase transition from N to FFLO SF phase, $E_{G}$ should decrease as the onset of $\Delta$. The minimization of $E_{G}$ with respect to $\Delta$ defines a function $S\left(t_{\textrm{AM}},{\bf q},\Delta\right)$:
@@
Equilibrium FFLO states satisfy $S\left(t_{\textrm{AM}},{\bf q},\Delta\right)=0$, which is also known as the gap equation. In the following, this gap equation is utilized to find the phase boundary between N and FFLO SF phase. We note that, $S(t_{\textrm{AM}},{\bf q},\Delta)$ can alternatively be understood as the inverse vertex function (i.e., the Cooper-pair propagator in the normal state with frequency $\omega=\Delta$) in the ladder approximation Sheehy2015 ; Hu2025-2 . Therefore, the condition $S(t_{\textrm{AM}},{\bf q},0)=0$ is precisely the well-known Thouless criterion for the onset of the N to SF phase transition.
@@ IV Phase diagram at ${\bf q}={\bf 0}$
In this section, we determine the phase boundaries assuming ${\bf q}={\bf 0}$ pairing, which amounts to neglecting the possibility of the FFLO state. Within this approximation, $A_{{\bf k}}=\xi_{{\bf k}}$ and $E_{{\bf k\pm}}=\sqrt{\xi_{{\bf k}}^{2}+\Delta^{2}}\pm J_{\mathbf{k}}$. In our terminology, $E_{{\bf k\pm}}$ is non-negative. If $E_{{\bf k\pm}}>0$ for all $\bf k$}, such a SF phase is called BCS SF. If $E_{{\bf k\pm}}=0$ at certain $\bf k$}, then the Bogoliubov quasiparticle has a “Fermi surface” structure, which is also known as BFS (Volovik1993, ). As discussed below, we have nodal SF with BFS protected by $\mathbb{Z}_{2}$ topological invariant. In the following, we first investigate the BCS SF phase.
@@ IV.1 BCS SF phase
$\begin{split}\frac{\Delta}{\varepsilon_{F}}&=\sqrt{\frac{\varepsilon_{B}}{\varepsilon_{F}}},\\ \frac{\mu}{\varepsilon_{F}}&=1-\frac{1}{2}\frac{\varepsilon_{B}}{\varepsilon_{F}}.\end{split}$ (26)
Here, we want to emphasize that $t_{\textrm{AM}}$ does not appear in Eq. (24), Eq. (25), and Eq. (26). This is in agreement with the earlier result in Ref. (Sheehy2015, ), where a small magnetic field does not appear in the stationary pairing amplitude of a BCS SF.
@@ IV.2 Fixed chemical potential
Next we integrate over the first term. Introducing polar coordinate ${\bf k}=[k\cos\left(\phi\right),k\sin\left(\phi\right)]$ and setting $E_{{\bf k\eta}}=0$, the boundary $k_{\pm}(\phi)$ is already shown in Fig. 1(a) with the following expression:
@@
$\int_{0}^{2\pi}\frac{1}{\frac{2-t_{\textrm{AM}}^{2}}{t_{\textrm{AM}}^{2}}+\cos(\phi)}d\phi=\frac{2\pi}{\sqrt{\left(\frac{2-t_{\textrm{AM}}^{2}}{t_{\textrm{AM}}^{2}}\right)^{2}-1}},$ (34)
and finally $I_{1}$ is expressed as:
@@
It is worth noting that the area of the ellipse in Fig. 1(a) is given by $\pi ab=\pi\mu/\sqrt{1-t_{\textrm{AM}}^{2}}$, therefore, the ground-state energy density of the N phase is proportional to the area of the ellipse. Such a geometric meaning can also be understood by setting $t_{\textrm{AM}}=0$, the above ground-state energy density becomes the well-known result $E_{G,\textrm{N}}/\mathcal{A}=-\mu^{2}/(4\pi)$.
@@
This determines the black solid curve in Fig. 2(a). When $t_{\textrm{AM}}$ is small, i.e. anisotropic spin splitting is weak, we have SF phase, even to the high chemical potential regime. Such a feature is in accordance with the earlier result that SF phase exist at high chemical potential regime when Zeeman field is small Sheehy2015 . Such SF becomes unstable when the anisotropic NRSS becomes large.
Figure 2: (a) Phase diagram at fixed chemical potential $\mu$ (normalized to the two-body binding energy $\varepsilon_{B}$) versus NRSS $t_{\textrm{AM}}$. (b) Phase diagram at fixed total particle number $N$, as a function of the binding energy $\varepsilon_{B}$ normalized to the Fermi energy $\varepsilon_{F}$. The black curves in both figures show the Pauli-limit, upper altermagnetic coupling strength, where a SF with zero center-of-mass momentum $\mathbf{q}=0$ becomes unstable towards a normal state, when the possibility of an inhomogeneous FFLO is not considered. In the weak-coupling regime (i.e., for large $\mu/\varepsilon_{B}$ or small $\varepsilon_{B}/\varepsilon_{F}$), those two black curves provide an excellent estimate for the phase boundary between a TBFS phase and a FFLO state, within the superfluid domain.
@@ IV.3 Fixed total particle number
After determining the phase boundary at fixed chemical potential, our next task is to consider the case of a fixed total particle number. By approaching the phase transition from the N phase, $\mu=\sqrt{1-t_{\textrm{AM}}^{2}}\varepsilon_{F}$ according to Eq. (36). Then Eq. (38) leads to
@@
which gives the black solid curve in Fig. 2(b) for $\varepsilon_{B}/\varepsilon_{F}<1/2$. When $t_{\textrm{AM}}=0$, arbitrarily weak attraction interactions are enough to drive the system into SF phase, as expected. The introduction of $t_{\textrm{AM}}$ will destabilize SF and the system transits to a normal state.
@@ IV.4 Nodal SF
Fig. 1(c) shows the typical behavior of $E_{\mathbf{k}}$ with respect to different $t_{\textrm{AM}}$ values: $E_{\mathbf{k}}$ acquires a global minimum at $\xi_{{\bf k}}=\frac{t_{\textrm{AM}}}{\sqrt{1-t_{\textrm{AM}}^{2}}}\Delta$ with the value $E_{\mathbf{k},\textrm{min}}=\frac{1-t_{\textrm{AM}}}{\sqrt{1-t_{\textrm{AM}}^{2}}}\Delta-\mu t_{\textrm{AM}}$.
BFS exists when $E_{\mathbf{k},\textrm{min}}=0$ has real positive root ($\bar{t}_{\textrm{AM}}$) within $[0,1)$. As $E_{\mathbf{k},\textrm{min}}=0$ is equivalent to the cubic equation
$t_{\textrm{AM}}^{3}+t_{\textrm{AM}}^{2}+\left(\frac{\Delta}{\mu}\right)^{2}t_{\textrm{AM}}-\left(\frac{\Delta}{\mu}\right)^{2}=0,$ (42)
@@
By inserting Eq. (24) into Eq. (42), we obtain the equation for determining $\bar{t}_{\textrm{AM}}$ at fixed $\mu$, which is displayed as the black dashed line in Fig. 2(a). At the same time, by plugging Eq. (26) in Eq. (42), we obtain the equation for determining $\bar{t}_{\textrm{AM}}$ at fixed total particle number, which is displayed as the black dashed line in Fig. 2(b). With $\bar{t}_{\textrm{AM}}<t_{\textrm{AM, c}}$, the nodal SF phase does exist in both fixed chemical potential and fixed total particle number cases. From Fig. 2(b), we know SF /nodal SF is preferred for a strong/weak pairing interactions, in accordance with Feiguin et al. (Feiguin2009, ).
@@ IV.5 Topological BFS
The fact that the BFS shown in Fig. 1(e)-(f) shares the same dimension as the normal state Fermi surface indicates it has a topological origin. To reveal the topological invariant, here we hinges on the Pfaffian Pf($\mathbf{k}$) of Eq. 7. Since Eq. 7 preserves both charge conjuration and parity symmetry, as suggested by Agterberg and co-workers (Agterberg2017, ), we can always find a suitable basis to transform Eq. (7) into a screw form whose Pf($\mathbf{k}$) is given by
@@
Here we would like to compare TBFS realized here with existing approaches. It is well known that BFS can not be realized in conventional 2D $s$-wave superconductors through external magnetic field alone, as the required magnetic field already drives the system into N phase before BFS can form. The replacement of the nodeless $s$-wave pairing by the nodal $d_{x^{2}-y^{2}}$-wave pairing has shown to enable TBFS (Yang1998, ; Setty2020-1, ). Alternatively, inducing supercurrent in $s$-wave superconductors with high Fermi velocity can also give rise to TBFS (Zhu2021, ). A further route involves multiband superconductors that preserve inversion symmetry but breaks time-reversal symmetry, as demonstrated in recent theoretical studies (Brydon2018, ; Lapp2020, ; Setty2020-2, ). In contrast to these approaches, AM introduces a new but simple mechanism for realizing TBFS. This scenario requires neither external magnetic field nor spin-orbit coupling, but only a single band and $s$-wave pairing, offering a minimal yet realistic platform for TBFS in superconducting systems.
@@ V FFLO phase boundary driven by NRSS
Now we turn to the FFLO phase. To study the FFLO phase, we assume a continuous transition, with decreasing $t_{\textrm{AM}}$, from the N phase to the FFLO phase, occurring when the curvature of $E_{G}$ versus $\Delta$, at $\Delta=0$, becomes 0 for ${\bf q}\neq(0,0)$. This equals to $S(t_{\textrm{AM}},{\bf q},0)=0$ with
@@
There is a duality between $I_{3}$ and $I_{4}$. By changing $t_{\textrm{AM}}$ to $t_{\textrm{AM}}$ and $\phi$ to $\phi+\frac{3\pi}{2}$ simultaneously, then $a,b$ becomes $b,a$ and $\sin(\phi),\cos(\phi)$ becomes $\cos(\phi),\sin(\phi)$, therefore $f_{2}^{+}(\phi)$ becomes $f_{2}^{-}(\phi)$, $f_{1}^{+}(\phi)$ becomes $f_{1}^{-}(\phi)$, and $f_{0}^{+}$ becomes $f_{0}^{-}$. Since the translation of integral variable $\phi\rightarrow\phi+\frac{3\pi}{2}$ does not change the integration over $\phi$, $I_{3}$ and $I_{4}$ should be connected by $t_{\textrm{AM}}\rightarrow-t_{\textrm{AM}}$. Such a fact simplifies the integrations. In the following, we firstly evaluate $I_{4}$ by contour integration, then changing $t_{\textrm{AM}}$ to $t_{\textrm{AM}}$ to obtain $I_{3}$.
@@
For $d_{xy}$-wave altermagnetic spin splitting, there are two characteristic high-symmetric directions for the FF momentum $\mathbf{q}$: along the $\left(1,0\right)$ direction (or $\left(0,1\right)$), and along the $\left(1,1\right)$ direction (or equivalently $\left(1,-1\right)$) according to the $C_{4z}$ rotational symmetry. In the following, we first analyze the case of the $\left(1,1\right)$ direction for concreteness, and then comment on the situation along the $\left(1,0\right)$ direction, which appears to be more difficult. As discussed in Appendix A, the angular dependence of the peak value of $S(t_{\textrm{AM}},{\bf q},0)/\mathcal{A}$ as a function of $q$ turns out to be weak. We thus anticipate that the analysis along the $\left(1,1\right)$ direction will provide a reasonably accurate phase boundary for the onset of the FFLO state.

### v2

@@ III.2 Thermodynamic potential and ground-state energy
At $T$ = 0 K, referred to as the ground-state energy, $\Omega$ takes the following form,
@@
where $\Theta(x)$ is the Heaviside step function. Since we are working on a continuum model, it is straightforward to convert the momentum sum to an integration, and the ground-state energy density becomes:
@@
where the renormalization condition Eq. (4) is used. At a given chemical potential, the phase boundary between SF and normal phase is obtained by equating the ground-state energies of these two phases:
@@ III.3 Gap equation
During the phase transition from normal to FFLO phase, $E_{G}$ should decrease as the onset of $\Delta$. The minimization of $E_{G}$ with respect to $\Delta$ defines a function $S\left(t_{\textrm{AM}},{\bf q},\Delta\right)$:
@@
Equilibrium FFLO states satisfy $S\left(t_{\textrm{AM}},{\bf q},\Delta\right)=0$, which is also known as the gap equation. In the following, this gap equation is utilized to find the phase boundary between normal and FFLO phase. We note that, $S(t_{\textrm{AM}},{\bf q},\Delta)$ can alternatively be understood as the inverse vertex function (i.e., the Cooper-pair propagator in the normal state with frequency $\omega=\Delta$) in the ladder approximation Sheehy2015 ; Hu2025-2 . Therefore, the condition $S(t_{\textrm{AM}},{\bf q},0)=0$ is precisely the well-known Thouless criterion for the onset of the normal to SF phase transition.
@@ IV Phase diagram at ${\bf q}={\bf 0}$
In this section, we determine the phase boundaries assuming ${\bf q}=0$ pairing, which amounts to neglecting the possibility of the FFLO state. Within this approximation, $A_{{\bf k}}=\xi_{{\bf k}}$ and $E_{{\bf k\pm}}=\sqrt{\xi_{{\bf k}}^{2}+\Delta^{2}}\pm J_{\mathbf{k}}$. In our terminology, $E_{{\bf k\pm}}$ is non-negative. If $E_{{\bf k\pm}}>0$ for all $\bf k$}, such a SF phase is called BCS SF. If $E_{{\bf k\pm}}=0$ at certain $\bf k$}, then the Bogoliubov quasiparticle has a ”Fermi surface” structure, which is also known as BFS (Volovik1993, ). As discussed below, we have nodal SF with BFS protected by $\mathbb{Z}_{2}$ topological invariant. In the following, we first investigate the BCS SF phase.
@@ IV.1 BCS SF phase
$\begin{split}\frac{\Delta}{\varepsilon_{F}}&=\sqrt{2\frac{\varepsilon_{B}}{\varepsilon_{F}}},\\ \frac{\mu}{\varepsilon_{F}}&=1-\frac{1}{2}\frac{\varepsilon_{B}}{\varepsilon_{F}}.\end{split}$ (26)
Here, we want to emphasize that $t_{\textrm{AM}}$ does not appear in Eq. (24), Eq. (25), and Eq. (26). This is in agreement with the earlier result in Ref. (Sheehy2015, ), where a small magnetic field does not appear in the stationary pairing amplitude of a BCS SF phase. Therefore, both altermagnetism and Zeeman field do not effect the BCS SF phase.
@@ IV.2 Fixed chemical potential
Next we integrate over the first term. Introducing polar coordinate ${\bf k}=[k\cos\left(\phi\right),k\sin\left(\phi\right)]$ where $k$ and $\phi$ are the radial distance and polar angle respectively, and setting $E_{{\bf k\eta}}=0$, the boundary $k_{\pm}(\phi)$ is already shown in Fig. 2(a) with the following expression:
@@
It is worth noting that the area of the ellipse in Fig. 2(a) is given by $\pi ab=\pi\mu/\sqrt{1-t_{\textrm{AM}}^{2}}$, therefore, the ground-state energy density of the normal phase is proportional to the area of the ellipse. Such a geometric meaning can also be understood by setting $t_{\textrm{AM}}=0$, the above ground-state energy density becomes the well-known result $E_{G,\textrm{N}}/\mathcal{A}=-\mu^{2}/(4\pi)$.
@@
This determines the black solid curve in Fig. 1(a). When $t_{\textrm{AM}}$ is small, i.e. anisotropic spin splitting is weak, we have BCS SF phase, even to the high chemical potential regime. Such a feature is in accordance with the earlier result that BCS SF phase exist at high chemical potential regime when Zeeman field is small Sheehy2015 . Large anisotropic NRSS will destabilize the BCS SF phase and drive a phase transition.
@@ IV.3 Fixed total particle number
After determining the phase boundary at fixed chemical potential, our next task is to consider the case of a fixed total particle number. By approaching the phase transition from the normal phase, $\mu=\sqrt{1-t_{\textrm{AM}}^{2}}\varepsilon_{F}$ according to Eq. (35). Then Eq. (37) leads to
@@
which gives the black solid curve in Fig. 1(b) for $\varepsilon_{B}/\varepsilon_{F}<1/2$. When $t_{\textrm{AM}}=0$, arbitrarily weak attraction interactions are enough to drive the system into the BCS SF phase, as expected. The introduction of $t_{\textrm{AM}}$ will destabilize BCS SF phase and the system transits to a normal state.
@@ IV.4 Nodal SF
Fig. 2(c) shows the typical behavior of $E_{\mathbf{k}}$ with respect to different $t_{\textrm{AM}}$ values: $E_{\mathbf{k}}$ acquires a global minimum at $\xi_{{\bf k}}=\frac{t_{\textrm{AM}}}{\sqrt{1-t_{\textrm{AM}}^{2}}}\Delta$ with the value $E_{\mathbf{k},\textrm{min}}=\sqrt{1-t_{\textrm{AM}}^{2}}\Delta-\mu t_{\textrm{AM}}$.
BFS exists when $E_{\mathbf{k},\textrm{min}}=0$ has real positive root ($\bar{t}_{\textrm{AM}}$) within $[0,1)$. Since $E_{\mathbf{k},\textrm{min}}=0$ leads to
$\bar{t}_{\textrm{AM}}=\sqrt{\frac{\Delta^{2}}{\Delta^{2}+\mu^{2}}},$ (41)
@@
By inserting Eq. (24) into Eq. (41), we obtain the equation for determining $\bar{t}_{\textrm{AM}}$ at fixed $\mu$:
$\bar{t}_{\textrm{AM}}=\frac{\sqrt{1+2\mu/\varepsilon_{B}}}{1+\mu/\varepsilon_{B}},$ (42)
which is displayed as the black dashed line in Fig. 1(a). At the same time, by plugging Eq. (26) in Eq. (41), we obtain the equation for determining $\bar{t}_{\textrm{AM}}$ at fixed total particle number:
$\bar{t}_{\textrm{AM}}=\frac{\sqrt{2\varepsilon_{B}/\varepsilon_{F}}}{1+\varepsilon_{B}/(2\varepsilon_{F})}.$ (43)
this curve is above the black line in Fig. 1(b). Therefore, the nodal SF phase exists in fixed chemical potential case as $\bar{t}_{\textrm{AM}}<t_{\textrm{AM, c}}$, but vanishes in fixed total particle number cases.
@@ IV.5 Topological BFS
The fact that the BFS shown in Fig. 2(e)-(f) shares the same dimension as the normal state Fermi surface indicates it has a topological origin. To reveal the topological invariant, here we hinges on the Pfaffian Pf($\mathbf{k}$) of Eq. (7). Since Eq. (7) preserves both charge conjuration and parity symmetry, as suggested by Agterberg and co-workers (Agterberg2017, ), we can always find a suitable basis to transform Eq. (7) into a screw form whose Pf($\mathbf{k}$) is given by
@@
Here we would like to compare TBFS realized here with existing approaches. It is well known that BFS can not be realized in conventional 2D $s$-wave superconductors through external magnetic field alone, as the required magnetic field already drives the system into normal phase before BFS can form. The replacement of the nodeless $s$-wave pairing by the nodal $d_{x^{2}-y^{2}}$-wave pairing has shown to enable TBFS (Yang1998, ; Setty2020-1, ). Alternatively, inducing supercurrent in $s$-wave superconductors with high Fermi velocity can also give rise to TBFS (Zhu2021, ). A further route involves multiband superconductors that preserve inversion symmetry but breaks time-reversal symmetry, as demonstrated in recent theoretical studies (Brydon2018, ; Lapp2020, ; Setty2020-2, ). In contrast to these approaches, AM introduces a new but simple mechanism for realizing TBFS. This scenario requires neither external magnetic field nor spin-orbit coupling, but only a single band and $s$-wave pairing, offering a minimal yet realistic platform for TBFS in superconducting systems.
@@ V FFLO phase boundary driven by NRSS
Now we turn to the FFLO phase. To study the FFLO state, we assume a continuous transition, with decreasing $t_{\textrm{AM}}$, from the normal to the FFLO phase, occurring when the curvature of $E_{G}$ versus $\Delta$, at $\Delta=0$, becomes 0 for ${\bf q}\neq(0,0)$. This equals to $S(t_{\textrm{AM}},{\bf q},0)=0$ with
@@
There is a duality between $I_{3}$ and $I_{4}$. By changing $t_{\textrm{AM}}$ to $t_{\textrm{AM}}$ and $\phi$ to $\phi+\frac{3\pi}{2}$ simultaneously, then $a,b$ becomes $b,a$ and $\sin(\phi),\cos(\phi)$ becomes $\cos(\phi),\sin(\phi)$, therefore $f_{2}^{+}(\phi)$ becomes $f_{2}^{-}(\phi)$, $f_{1}^{+}(\phi)$ becomes $f_{1}^{-}(\phi)$, and $f_{0}^{+}$ becomes $f_{0}^{-}$. Since the translation of integral variable $\phi\rightarrow\phi+\frac{3\pi}{2}$ does not change the integration over $\phi$, $I_{3}$ and $I_{4}$ should be connected by $t_{\textrm{AM}}\rightarrow-t_{\textrm{AM}}$. Such a fact simplifies the integrations. For example, we can firstly evaluate $I_{4}$ by contour integration, then changing $t_{\textrm{AM}}$ to $t_{\textrm{AM}}$ to obtain $I_{3}$.
@@
For $d_{xy}$-wave altermagnetic spin splitting, there are two high-symmetric directions for the FF momentum $\mathbf{q}$: along the $\left(1,0\right)$ direction (or $\left(0,1\right)$), and along the $\left(1,1\right)$ direction (or equivalently $\left(1,-1\right)$) according to the $C_{4z}$ rotational symmetry. The analytic evaluation of $I_{3}$ and $I_{4}$ is highly nontrivial; the calculation for $\bf{q}$ along $\left(1,1\right)$ direction in presented in Appendix A. By contrast, the corresponding integrals for $\bf{q}$ along $\left(1,0\right)$ direction are considerably more challenging, and we have not found a closed-form expression. As discussed in Appendix B, the angular dependence of the peak value of $S(t_{\textrm{AM}},{\bf q},0)/\mathcal{A}$ as a function of $q$ turns out to be weak. We thus anticipate that the analysis along the $\left(1,1\right)$ direction will provide a reasonably accurate phase boundary for the onset of the FFLO state.
Figure 3: Plot of $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$ (see Eq. (50) for definition) at different $t_{\textrm{AM}}$ values, here $\varepsilon_{B}=0.05$ and $\mu=1$.
V.1 Fixed chemical potential
With $I_{2}$, $I_{3}$, and $I_{4}$ evaluated, introducing the dimensionless momentum $Q\equiv q/\sqrt{\mu}$ and $Q_{c}\equiv\sqrt{\frac{1}{1-t_{\textrm{AM}}}}-\sqrt{\frac{1}{1+t_{\textrm{AM}}}}$, $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$ now is calculated as:
$\frac{S(t_{\textrm{AM}},Q,0)}{\mathcal{A}}=\frac{1}{4\pi}\begin{cases}\ln\sqrt{\frac{2\varepsilon_{B}}{\mu}}-\ln\left|\frac{a-b}{a+b}\right|-\ln\left|\sqrt{4-(1-t_{\textrm{AM}}^{2})Q^{2}}\right|,Q\in(0,Q_{c}]\\ \ln\sqrt{\frac{2\varepsilon_{B}}{\mu}}-\ln\left|\frac{a-b}{a+b}\right|+\ln\left|\frac{2t_{\textrm{AM}}}{\sqrt{C(1)}+(1-t_{\textrm{AM}}^{2})+\sqrt{2D_{+}(1)}}\right|,Q\in(Q_{c},\sqrt{\frac{1}{1-t_{\textrm{AM}}}}+\sqrt{\frac{1}{1+t_{\textrm{AM}}}}]\end{cases}$ (50)
where $C(1)=(1-t^{2}_{AM})^{2}Q^{2}+8t^{2}_{AM}$ and $D_{+}(1)=(1-t^{4}_{AM})Q^{2}+(1-t^{2}_{AM})Q\sqrt{C(1)}-4t^{2}_{AM}$.
Fig. 3 plot $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$ for $\varepsilon_{B}=0.05$ and $\mu=1$. From which, it is clear that $S(t_{\textrm{AM}},Q,0)$ increases slowly for $Q<Q_{c}$ and decreases rapidly for $Q>Q_{c}$, therefore $Q_{c}$ is a kink for $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$ which reaches the maximum value:
It is well known that the normal phase is stable at $S(t_{\textrm{AM}},Q,0)/\mathcal{A}<0$, therefore we see that the FFLO phase emerges at $Q=Q_{c}$ and $S(t_{\textrm{AM}},Q,0)/\mathcal{A}=0$, the simultaneous solution of which leads to the critical $t_{\textrm{AM}}$ for FFLO phase:
In Fig. 1(a), we plot the FFLO phase boundary $t_{\textrm{AM, FFLO}}$ as the red curve. When $\mu<\varepsilon_{B}$, $S(t_{\textrm{AM}},Q_{c},0)=0$ has no solution and the system remains in the normal phase. For $\mu\geq\varepsilon_{B}$ the FFLO phase emerges from the normal phase side of the phase diagram, in agreement with our assumption that the decrease of $t_{\textrm{AM}}$ leads to a continuous normal-FFLO phase transition. From Fig. 1(a), $t_{\textrm{AM}}$ has the largest window for observing FFLO phase when $\mu=\varepsilon_{B}$. After that, the window gradually shrinks.
In Soto-Garrido et al. (Soto-Garrido2014, ), the integration over $\bf{k}$ is carried out within a shell around the Fermi surface, and such a finite range of integration leads to a flat $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$. Based on such a result, Soto-Garrido et al. (Soto-Garrido2014, ) claimed the absence of a FFLO phase. Our results reveal the importance of full $\mathbf{k}$-integration in restoring the FFLO phase.
V.2 Fixed total particle number
Figure 4: (a) Illustration of Fermi surface nesting for $\mathbf{q}$ along (1, 1) direction; (b) Bogoliubov Fermi surfaces for $\mathbf{q}=(q_{c}/\sqrt{2},q_{c}/\sqrt{2})$, here $q_{c}=a-b$. (c) Fermi surface nesting for $\mathbf{q}$ along (1, 0) direction. (d) Bogoliubov Fermi surfaces for $\mathbf{q}=(q_{c},0)$, here $q_{c}$ is given by Eq. (54). $\Delta=0$, $\mu=1$, and $t_{AM}=0.5$ are chosen for plotting Bogoliubov Fermi surfaces in (b) and (d).
V.3 Geometric interpretation
In previous section, we have shown that, for $\Delta=0$, $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$ exhibits a nonanalytic dependence on $Q$ according to the contour integration of complex analysis. The nonanalyticity occurs at $Q_{c}=\sqrt{\frac{1}{1-t_{\textrm{AM}}}}-\sqrt{\frac{1}{1+t_{\textrm{AM}}}}$ which corresponds to $q_{c}=a-b$. Such an expression of $q_{c}$ is in agreement with Zhang et al. (ZhangS2024, ) where the high chemical potential approximation is made. Such a critical momentum admits a clearly geometrical interpretation as shown in Fig. 4(a), $q_{c}$ is nothing but the momentum shift to make two of the four crossings between the two Fermi ellipses ($E_{{\bf k}\pm}=A({\bf k})\pm B({\bf k})=0$) and the black dashed circle ($A({\bf k})=0$) become one tangent point.
The obtained BFS is shown in Fig. 4(d), where the number of BFS decreases from 4 to 2. As elaborated in Appendix B, our numerical simulation verifies that Eq. (54) indeed gives a good estimate of the actual $q_{c}$ when $\mathbf{q}$ is along the $(1,0)$ direction.
VI CONCLUSIONS AND OUTLOOKS
In summary, we have analytically investigated the ground-state phase diagram of 2D spin-1/2 Fermi gases exhibiting $d$-wave altermagnetic spin splitting in the presence of $s$-wave pairing, within the framework of BdG mean-field theory. Under both fixed chemical potential and fixed total particle number conditions, our phase diagram illustrated in Fig. 1 definitely confirms the existence of FFLO states under $s$-wave pairing, thereby resolving previous contradictory conclusions in the literatures. In addition, we provide a geometric interpretation of the onset $\mathbf{q}_{c}$ of FFLO states, offering intuitive insight into its origin. Quite unexpectedly, we also discover a nodal SF, characterized by the TBFS at fixed chemical potential. Given the simplicity of the system - a single band, without external magnetic field, and under nodeless $s$-wave pairing - the presence of TBFS is both surprising and compelling.
Fig. 1 illustrates the fundamental relationship between anisotropic NRSS and attractive pairing interaction. When the attractive pairing interaction is strong, BCS SF phase is expected with a zero center-of-mass momentum and the Fermi surfaces are fully gapped. Meanwhile, for a very large anisotropic NRSS, the metallic normal phase is anticipated. Nodal SF and FFLO occur when anisotropic NRSS and attractive pairing are both at intermediate strength. All phases can be realized in either ultracold gases or electronic superconductors by choosing an appropriate $t_{\textrm{AM}}$. In the context of ultracold gases, a large $t_{\textrm{AM}}$ is required as shown in Fig. 1(a), so we can access to all four phases via tuning the two-body interactions $\varepsilon_{B}$ through Feshbach resonance Chin2010 . For electronic superconductors, a small $t_{\textrm{AM}}$ suffices, as shown in Fig. 1(b). By increasing the Fermi energy $\varepsilon_{F}$ via techniques like electrostatic gating Wu2023 , we can sequentially observe BCS SF, FFLO, and normal phase.
We conclude this section by noting several natural extensions of this work. First, incorporating additional physical ingredients - such as external magnetic fields, spin-orbit coupling, $d$-wave pairing, and finite temperature - could significantly enrich the phase diagram and reveal new emergent pairing phenomena. A detailed analysis of the effects of disorder will be presented in a separated work. Second, while our analysis is based on the simplest single-plane-wave ansatz here, higher plane-wave FFLO ansatz such as the LO state with $\Delta(\mathbf{x})=\Delta\cos(\mathbf{q}\cdot\mathbf{x})$ are expected to be energetically more favorable. Third, our analysis relies on the mean-field framework, which provides qualitatively robust results. Quantitative refinements, with the inclusion of quantum pair fluctuations (He2015, ) will be pursued in future work. Finally, the topological BFS uncovered in this work may provide a fertile platform for realizing Majorana bound states, opening the door to potential applications in topological quantum computation.
Acknowledgements.
Appendix A Integration of $I_{4}$
In this section, we demonstrate the evaluate of the integral $I_{4}$. There are two integral variables $k$ and $\phi$ in $I_{4}$. We can firstly integrate $k$ and then $\phi$ as the case in evaluating $I_{1}$ and $I_{2}$. Or we can integrate $\phi$ first and then $k$. It is found that the second choice is more convenient and is adopted here.